# An Expert is Worth One Token: Synergizing Multiple Expert LLMs as Generalist via Expert Token Routing

### v2

@@ Abstract
We present Expert-Token-Routing, a unified generalist framework that facilitates seamless integration of multiple expert LLMs. Our framework represents expert LLMs as special expert tokens within the vocabulary of a meta LLM. The meta LLM can route to an expert LLM like generating new tokens. Expert-Token-Routing not only supports learning the implicit expertise of expert LLMs from existing instruction dataset but also allows for dynamic extension of new expert LLMs in a plug-and-play manner. It also conceals the detailed collaboration process from the user’s perspective, facilitating interaction as though it were a singular LLM. Our framework outperforms various existing multi-LLM collaboration paradigms across benchmarks that incorporate six diverse expert domains, demonstrating effectiveness and robustness in building generalist LLM system via synergizing multiple expert LLMs.  [^1] Codes and datasets are available at https://github.com/godcherry/ExpertTokenRouting.
@@
Ziwei Chai1     Guoyin Wang2      Jing Su2      Tianjie Zhang1      Xuanwen Huang1      Xuwu Wang2      Jingjing Xu2      Jianbo Yuan2      Hongxia Yang2      Fei Wu1      Yang Yang1 1Zhejiang University      2ByteDance Inc zwchai@zju.edu.cn
@@ 1 Introduction
Large language models (LLMs)—notably, GPT-4 and LLaMA (Touvron et al., 2023)—have demonstrated remarkable capabilities across a wide spectrum of tasks. However, their performance and reliability in certain specialized domains sometimes still fall short of expectations (Wang et al., 2023). Motivated by this, the recent year has seen emergence in the development of "expert" LLMs, with a particular focus on tailoring a general LLM to excel in specific domains or tasks. For example, there are efforts to refine the pretraining corpus to develop expert LLMs with enhanced capabilities in specific domains, such as CodeLlama (Rozière et al., 2024) and DeepSeekMath (Shao et al., 2024). Furthermore, applying continual training or supervised fine-tuning (SFT) on specially selected corpora can lead to significant enhancements on targeted tasks like data analysis (Li et al., 2023b; Hu et al., 2024) and financial analysis (Chen et al., 2023), making a general LLM into a specialized expert.
@@ 2.2 Learning Expert Embeddings via Expert Query Set
This embedding matrix $W_{E}$ are the only tunable parameters. Therefore, similar to pioneering work on expanding the language modeling head of LLMs (such as ToolkenGPT (Hao et al., 2023)), training ETR doesn’t require gradient propagation through the main part of the LLM parameters, leading to training that is both more stable and efficient than prompt tuning (Lester et al., 2021) or prefix tuning (Li and Liang, 2021). Thus, the GPU memory usage for expert token embedding tuning is nearly the same to that of LLM inference.
@@ 3.1 Experimental Settings
We compare our framework with two types of baseline methods: the prompting-based expert collaboration methods and the router-based expert routing methods. It’s worthy noting that in the existing prompting-based methods, as exemplified by Meta-Prompting (Suzgun and Kalai, 2024), there are two distinct components in its instruction: collaborating with experts and forming expert identity, both of which are crucial to the effectiveness of this method. However, in this paper’s setting, an external expert LLM library are provided, removing the need to create experts via prompting. Therefore, we adapt the segment of the Meta-prompting instruction dedicated to expert LLM collaboration to serve as a comparative method, which we refer to as "Meta-prompting-E" in this paper. We also compare the multi-LLM debate (Du et al., 2023) method, as a specific case within prompting-based methods, where all the experts are invoked for all queries. In router-based methods, we compare with LLM-Blender (Jiang et al., 2023), which ranks the responses from each LLM. And we select the top-ranked result as the routing expert. In the original design of the LLM-Blender, it is also capable of fusing the top-k responses. This paper marks these two methods as LLM-Blender (Top 1) and LLM-Blender (Fused), respectively. It is noteworthy that Zooter (Lu et al., 2023) is also a high-related router-based method. We plan to compare with it in the future since its implementation hasn’t been released.
@@ 3.4 Expert Routing Accuracy
In this section, we compare the expert routing accuracy in Table 3. For example, in the MMLU-Expert dataset, the questions in electrical engineering category has an oracle expert LLM, which is fine-tuned on the synthetic electrical engineering knowledge dataset (Section 3.1). The accuracy is 100% when all the queries are routed to its corresponding domain expert LLM, where random routing has an accuracy of 16.67%. It can be seen that the LLM-Blender, of which the ranking model is trained based on rewards aligned with human preferences, doesn’t perform well in accurately routing to the domain expert. Meta-Prompting-E achieves a 67.08% expert routing accuracy on the MMLU-Expert dataset, while ETR reaches 82.11%, exceeding it by 15.03%.
@@ 4 Related Works
A bunch of methods aim at enhancing the reasoning capabilities of LLMs in specific tasks by instructing LLMs into experts through the crafting of prompts. These methods mainly focus on boosting a single LLM by simulating multiple expert LLM instances, and can also be applied to our paper’s setting. ExpertPrompting (Xu et al., 2023) elicits LLMs to generate expert-level responses using proper crafting of prompts. Meta-Prompting (Suzgun and Kalai, 2024) breaks down complex tasks into subtasks handled by expert instances of a single LLM. Multi-LLM Debate (Du et al., 2023) involves the use of multiple LLM instances to propose and debate their individual responses and reasoning processes over several rounds in order to converge on a common final answer.
@@ B.2 Expert Token Training Setup
We provide the hyper-parameters on expert token training in Table 7 for reproduction. For the sake of training stability, the initialization of the expert token is conducted using the mean value of embeddings found within the language modeling head of the meta LLM. It can be seen that when using Qwen-7B-Chat as the meta LLM, the parameters that need to be trained amount to only 24,567, which accounts for merely $3.51×10^{-6}$ of the total parameters.

### v3

@@ Abstract
We present Expert-Token-Routing, a unified generalist framework that facilitates seamless integration of multiple expert LLMs. Our framework represents expert LLMs as special expert tokens within the vocabulary of a meta LLM. The meta LLM can route to an expert LLM like generating new tokens. Expert-Token-Routing not only supports learning the implicit expertise of expert LLMs from existing instruction dataset but also allows for dynamic extension of new expert LLMs in a plug-and-play manner. It also conceals the detailed collaboration process from the user’s perspective, facilitating interaction as though it were a singular LLM. Our framework outperforms various existing multi-LLM collaboration paradigms across benchmarks that incorporate six diverse expert domains, demonstrating effectiveness and robustness in building generalist LLM system via synergizing multiple expert LLMs.  [^1] Codes are available at https://github.com/zjunet/ETR.
@@
Ziwei Chai1,2, Guoyin Wang3, Jing Su3, Tianjie Zhang1, Xuanwen Huang1, Xuwu Wang3 Jingjing Xu3, Jianbo Yuan3, Hongxia Yang3, Fei Wu1, Yang Yang1thanks:   Corresponding authors. 1 Zhejiang University, 2 Beijing Huairou Laboratory, 3 ByteDance Inc. {zwchai, yangya}@zju.edu.cn
@@ 1 Introduction
Large language models (LLMs)—notably, GPT-4 and LLaMA (Touvron et al., 2023)—have demonstrated remarkable capabilities across a wide spectrum of tasks. However, their performance and reliability in certain specialized domains sometimes still fall short of expectations (Wang et al., 2023; Zhao et al., 2024). Motivated by this, the recent year has seen emergence in the development of "expert" LLMs, with a particular focus on tailoring a general LLM to excel in specific domains or tasks. For example, there are efforts to refine the pretraining corpus to develop expert LLMs with enhanced capabilities in specific domains, such as CodeLlama (Rozière et al., 2024) and DeepSeekMath (Shao et al., 2024). Furthermore, applying continual training or supervised fine-tuning (SFT) on specially selected corpora can lead to significant enhancements on targeted tasks like data analysis (Li et al., 2023b; Hu et al., 2024) and financial analysis (Chen et al., 2023), making a general LLM into a specialized expert.
@@ 2.2 Learning Expert Embeddings via Expert Query Set
This embedding matrix $W_{E}$ are the only tunable parameters. Therefore, similar to pioneering work on expanding the language modeling head of LLMs (such as ToolkenGPT (Hao et al., 2023)), training ETR does not require gradient propagation through the main part of the LLM parameters, leading to training that is both more stable and efficient than prompt tuning (Lester et al., 2021) or prefix tuning (Li and Liang, 2021). Thus, the GPU memory usage for expert token embedding tuning is nearly the same to that of LLM inference.
@@ 3.1 Experimental Settings
We compare our framework with two types of baseline methods: the prompting-based expert collaboration methods and the router-based expert routing methods. It’s worthy noting that in the existing prompting-based methods, as exemplified by Meta-Prompting (Suzgun and Kalai, 2024), there are two distinct components in its instruction: collaborating with experts and forming expert identity, both of which are crucial to the effectiveness of this method. However, in this paper’s setting, an external expert LLM library are provided, removing the need to create experts via prompting. Therefore, we adapt the segment of the Meta-prompting instruction dedicated to expert LLM collaboration to serve as a comparative method, which we refer to as "Meta-prompting-E" in this paper. We also compare the multi-LLM debate (Du et al., 2023) method, as a specific case within prompting-based methods, where all the experts are invoked for all queries. In router-based methods, we compare with LLM-Blender (Jiang et al., 2023), which ranks the responses from each LLM. And we select the top-ranked result as the routing expert. In the original design of the LLM-Blender, it is also capable of fusing the top-k responses. This paper marks these two methods as LLM-Blender (Top 1) and LLM-Blender (Fused), respectively. It is noteworthy that Zooter (Lu et al., 2023) is also a high-related router-based method. We plan to compare with it in the future since its implementation has not been released.
@@ 3.4 Expert Routing Accuracy
In this section, we compare the expert routing accuracy in Table 3. For example, in the MMLU-Expert dataset, the questions in electrical engineering category has an oracle expert LLM, which is fine-tuned on the synthetic electrical engineering knowledge dataset (Section 3.1). The accuracy is 100% when all the queries are routed to its corresponding domain expert LLM, where random routing has an accuracy of 16.67%. It can be seen that the LLM-Blender, of which the ranking model is trained based on rewards aligned with human preferences, does not perform well in accurately routing to the domain expert. Meta-Prompting-E achieves a 67.08% expert routing accuracy on the MMLU-Expert dataset, while ETR reaches 82.11%, exceeding it by 15.03%.
@@ 3.7 Running Case Analysis
3.8 Running Time Analysis
We evaluate the running time of ETR in two different conditions: no active LLM switching (line 9 in Algorithm 1) versus active LLM switching (line 11 in Algorithm Algorithm 1). We randomly select 100 queries to calculate the average running time. To eliminate the impact of different response length by the two conditions on the running time, the maximum generation length is set on the shorter response length of the two (greedy decoding is used for deterministic response). The experimental results are as shown in Table 5.
Table 5: Time Cost of LLM Switching
Running Time (s)
No Active LLM Switching 1.589
Active LLM Switching 1.616
The experiment runs with CUDA 12.1, Pytorch 2.1 and Flash-Attention 2. Each LLM is served on an A100-SXM4-80G GPU, utilizing bf16 quantization. We notice that active LLM switching results in roughly 1.7% extra running time costs. This ratio is expected to further decrease as the generated response becomes longer.
@@ 4 Related Works
A bunch of methods aim at enhancing the reasoning capabilities of LLMs in specific tasks by instructing LLMs into experts through the crafting of prompts. These methods mainly focus on boosting a single LLM by simulating multiple expert LLM instances, which can also be applied to the setting in this paper. ExpertPrompting (Xu et al., 2023) elicits LLMs to generate expert-level responses using proper crafting of prompts. Meta-Prompting (Suzgun and Kalai, 2024) breaks down complex tasks into subtasks handled by expert instances of a single LLM. Multi-LLM Debate (Du et al., 2023) involves the use of multiple LLM instances to propose and debate their individual responses and reasoning processes over several rounds in order to converge on a common final answer.
@@ 6 Limitations
Ethics Statement
This work has been conducted using open-source Qwen Bai et al. (2023) models and aims to study methods for collaborating multiple expert language models within a singular generalist framework. We acknowledge that language models can exhibit biases and generate inappropriate content. The expert models were created using synthetic data generated by commercial language models, and data filtering was performed to avoid harmful content. Thus, we believe our work does not pose severe ethical concerns.
Acknowledgment
This work is supported by National Natural Science Foundation of China (No. 62322606, No. 62441605) and SMP-IDATA Open Youth Fund.
@@ B.2 Expert Token Training Setup
We provide the hyper-parameters on expert token training in Table 8 for reproduction. For the sake of training stability, the initialization of the expert token is conducted using the mean value of embeddings found within the language modeling head of the meta LLM. It can be seen that when using Qwen-7B-Chat as the meta LLM, the parameters that need to be trained amount to only 24,567, which accounts for merely $3.51\times 10^{-6}$ of the total parameters.